Solve for x present your answer in decimal form
- 4.5x + 2.4 = - 3.7x - 3.1

Answer by Alan3354(69443)

Solve for x present your answer in decimal form
- 4.5x + 2.4 = - 3.7x - 3.1
It's one step at a time.
Add 3.1
-4.5x + 5.5 = -3.7x
Add 4.5x
5.5 = 0.8x
Divide by 0.8
x = 5.5/0.8 = 55/8
x = 6.875